OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

i know its alot but please help

OpenStudy (anonymous):

it would be a lot easier to understand if you used the equation editor it is not that it is a lot to read, it is that it is unclear

OpenStudy (anonymous):

\[3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}\] is what i am reading it is close?

OpenStudy (anonymous):

it says math processing error

OpenStudy (anonymous):

refresh your browser